# State-Space Modeling in Long Sequence Processing: A Survey on Recurrence in the Transformer Era

### v1

@@ 3.3 Modeling Recurrence
where $\text{sg}(\cdot)$ denotes an operation (i.e., stop-gradient) which avoids the gradient flow during Backpropagation, [|][$\cdot$|$\cdot$][ ⋅ | ⋅ ] denotes concatenation and $\text{tr}^{l}(\cdot)$ the $l$-th Transformer layer. Notice that the recurrent dependency shifts one-layer downward per segment, as depicted in Figure 8, which differs from the same-layer recurrence in conventional RNNs. Thus, the largest possible dependency length grows linearly w.r.t. the number of layers as well as the segment length. Additionally, it is possible to not limit the state-cache solely to the previous segment, by storing the last $m$ states and using them as the extra context when processing the current segment. Thus, a Transformer-XL with $N$ layers and with a memory of size $m$, has a maximum temporal range of $N\times m$ with a memory complexity in the attention operation of $\mathcal{O}(T^{2}+Tm)$ when processing a segment with length $T$.
@@
In RMT the Effective context length for recurrence with memory is not limited by the depth of the network, which is the case for the cache of Transformer-XL. Moreover, while Transformer-XL stores $m\times T$ vectors per segment, RMT stores only $m$ memory vectors per segment. RMT is trained with BPTT, and the number of segments to backpropagate is a hyperparameter of the training procedure (the authors tested from 0 to 4 segments). Differently from Transformer-XL, during the backward pass memory-related gradients are not stopped between segments. Figure 8 depicts the architectural differences between Transformer-XL and RMT. In a recent technical report [29], the authors leverage the RMT architecture to extend the context length of a BERT model [139] up to two million tokens, while maintaining high memory retrieval accuracy. Another approach that still falls in the category of models described so far is Infini-Transformer [201], which incorporates a compressive memory into a vanilla attention mechanism. It also builds in both causal local attention and long-term linear attention mechanisms into a single Transformer block. Inspired by Dai et al. [51], the authors remark how performing attention solely on the current segment (that can be seen as a form of local attention) discards the attention states of the previous one. To counteract this issue, they propose to maintain the entire context history in the form of a compressive memory. An Infini-Transformer layer contains both global compressive and local fine-grained states. In fact, it maintains as many parallel compressive memories as the number of attention heads, in addition to the dot-product attention. Each segment is processed via a classic softmax self-attention, producing an output $Y_{\text{loc}}$. Then, a liner attention (see Eq. (9)) is exploited to retrieve content from the previous segment memory/state, $M_{s-1}$, as follows,
@@
where $\phi(x)=\text{elu}(x)+1$, $s$ is an index over the segments and $z_{.}$ is the normalization factor of linear attention (see Eq. (9), here we used the parallel form, in matrix notation). State/memory update is implemented by a delta-rule like mechanisms [259] (see Section 3.1), first retrieving existing entries (values) and subtracting them from the new values, before applying the associative bindings, as follows,
@@
Chunk-level Recurrences.   The just introduced segment-based approach allows networks to process very long texts, sequentially, keeping a recurrent memory/context. Another category of emerging approaches of recurrent models that handle portions of text, consists in processing sub-portions of the input sequence, referred to as chunks, by dividing the sequence into non-overlapping chunks and performing (serial) inter-chunk recurrent computations followed by (parallel) intra-chunk computations. Such a chunk-wise parallel form yields sub-quadratic complexity. Temporal Latent Bottleneck (TLB) [58] divides the input sequence $U=(u_{1},\dots,u_{L})$ into chunks of fixed size $C$, resulting in $\lfloor L/C\rfloor$ chunks that are sequentially processed one after the other. We denote with $U_{[i]}$ the $i$-th chunk of the input sequence, i.e., $U_{[i]}\vcentcolon=U_{iC:(i+1)C}\in\mathbb{R}^{C\times{d_{\text{in}}}}$. Each chunk is processed by a fast-stream (also referred to as perceptual module), implemented by a Transformer $\text{tr}(\cdot)$. The fast-stream computation is conditioned via cross-attention on information coming from a slow-stream, referred to as Temporal Latent Bottleneck $\mathcal{G}$, that aggregates cues across chunks and is updated once per chunk. Such a slow stream is computed recurrently, and it manages a state ${X}$ composed of a set of $N$ ${d_{\text{s}}}$-dimensional vectors. The state update in $\mathcal{G}(\cdot)$ is performed via a cross attention operation in which the queries are obtained projecting $X$ while the keys and values come from the output of the perceptual module. Overall, the model performs the following operation,
@@
The recurrent update of the TLB state $X$ is performed at lower rates with respect to the computations happening in the perceptual module, fostering the distillation of more stable, condensed and slowly changing features, while the perceptual module captures faster changing local cues. From the computational point of view, leveraging a chunk-based approach allows to achieve a complexity of $\mathcal{O}\left(\frac{L}{C}(C^{2}d+CN)\right)$, where $N$ is the number of temporal latent bottleneck state vectors. Hence, it has a much lower computational complexity compared to a vanila Transformer applied on the overall sequence. A concurrent work, Block-Recurrent Transformer (Brect) [114], handles the input sequence in a similar manner with respect to Transformer-XL: a document is split into multiple segments of size $T$, processed one after the other. Each segment is processed in chunks (or blocks) using a sliding window attention with size $C$, with a causal mask that forces each token to attend solely to the previous ones. Brect is composed by a recurrent cell that receives as inputs $C$ token embeddings, where $C$ is the block/window size, and a set of $N$ state vectors. Similarly to TLB, a two-way processing is performed in the proposed recurrent cell, consisting of the so-called vertical and horizontal “directions”: the vertical direction (i.e., corresponding to the fast-stream in TLB) is implemented by a Transformer layer that performs self-attention over the input tokens of the current block and cross-attends to the recurrent states, producing output token embeddings for the current block. The horizontal direction (i.e., the slow-stream in TLB) performs self-attention over the current state vectors, and cross-attends to the block input tokens, producing the next state vectors. Differently from TLB, cross attention and self attention are computed in parallel. Recurrence is integrated with the sliding window attention mechanism, since keys and values from the previous window are stored into a differentiable cache and concatenated to the ones of the current block. Residual connections are substituted by gating mechanisms (implemented either as fixed convex combinations or trainable LSTM-like gates) that help the model in forgetting redundant information. For every layer of the architecture, the last recurrent states of a segment are stored in a non-differentiable cache and fed to the following segment in the document as a warm-start. This mechanism extends the sliding window to cover the entire sequence. The cache implements a form of truncated BPTT over long documents. Block-State Transformer [225] substitutes the recurrent cell in Brect with a linear state-space model (see Section 4), which processes the sequence and preserves long-range dependencies, allowing also for parallel computations. Recurrent Attention Networks (RAN)[158] iteratively process the input sequence by means of non-overlapping windows, each of them processed via multi-head self-attention. Intra-window information is propagated by a global perception cell (GPC) vector, extracted from the self-attention representation of the current window. The GPC vector is concatenated with the input tokens of the next window. A memory review mechanism cross-attends the concatenation of the GPC vectors of all the windows to build the final output, encoding both contextual and global information. We report in Figure 9 the main differences in the processing schemes of segment level and chunk-level approaches, reporting examples of some of the described models.
@@ 3.4 Decaying Linear Attention by Gating Functions
Gating mechanisms were introduced since the dawn of RNNs, being them a key feature of the popular LSTM model [105, 235]. In previous subsections and in the context of linear approximations to attention mechanisms, we described several works connecting Linear Transformers to Fast weight Programmers (FWP) which introduced decay/gating mechanisms inherited from FWP [259, 118, 119, 179]. The original gating functions in LSTMs consisted of units with sigmoidal activations, that are responsible of gating specific signals (input, output, state) by means of multiplication. Lately, the concept of gating has been relaxed to consider any multiplicative interaction possibly followed by an activation function (e.g., element-wise multiplicative components that do not interact along the sequence length are referred to as gating mechanisms [110, 189]). Nevertheless, the role of gating has become increasingly popular and it is nowadays pivotal in many works.
Multiplicative Interactions.   Hua et al. [110] remarked that despite the linear theoretical complexity claimed by linear attention methods [134], they have not been able to supersede vanilla Transformers as the dominant choice in state-of-the-art systems. They attributed this to ($i$) approximations needed to achieve efficiency; ($ii$) non-trivial gap between theoretical complexity and empirical speed on accelerators, due to memory re-formatting and operations (see also Section 8.2); ($iii$) slow training on causal autoregressive tasks, due to sequential processing of the recurrent forms. To counteract these issues, the authors proposed FLASH, which is based on a novel layer, dubbed Gated Attention Unit (GAU), to substitute softmax multi-head self-attention. The key idea is to combine a Gated Linear Unit (GLU) [^9] A GLU [56] is an MLP which output is modulated via a gating (i.e., a learned multiplicative interaction). In fact, the layer input $U$ is projected by two learnable matrices $W_{p}\in\mathbb{R}^{{d_{\text{in}}}\times d_{e}}$ and $W_{t}\in\mathbb{R}^{{d_{\text{in}}}\times d_{e}}$ into two representations $P,T\in\mathbb{R}^{L\times d_{e}}$, that interact in an element-wise multiplicative manner to produce the layer output, as follows: $R=\sigma(UW_{p})\quad V=\sigma(UW_{t})\quad Y=(R\odot V)W_{y}.$ [56] and attention in a unified layer, as depicted in Figure 10.
@@
where $W_{r},W_{v}\in\mathbb{R}^{{d_{\text{in}}}\times d_{e}}$, $W_{z}\in\mathbb{R}^{{d_{\text{in}}}\times d_{z}}$ with $d_{z}\ll{d_{\text{k}}}$, $\mathcal{Q},\mathcal{K}$ are learned transformations that apply per-dim scalars and offsets to $Z$, $W_{y}\in\mathbb{R}^{d_{e}\times{d_{y}}}$ denotes an output learnable matrix, and $b_{v}\in\mathbb{R}^{d_{z}}$ are relative position bias. Notice that this formulation substitutes the softmax with the squared ReLU[272] (denoted with $\text{ReLU}^{2}$). This single-headed softmax-free formulation is capable to achieve multi-head Transformers performances without quality loss, and is cheaper in terms of computational requirements, by adding solely the learnable matrix $W_{z}$ with ${d_{\text{in}}}\times z$ parameters on top of the GLU. Additionally, the authors analyze the causes underlying the aforementioned memory and speed issues (issues $ii$ and $iii$). In fact, despite the huge advantages brought by the constant-inference computation and memory cost in autoregressive decoding (i.e., thanks to the stateful representation $S_{t}$) the rearrangement of computations in linear attention lead to an inefficiency in the case of autoregressive training. [^10] Due to the causal constraint for auto-regressive training, the query vector corresponds to a different cache value at each time step. This requires the model to compute and cache $L$ different values of the incremental state and requires $L$ memory accesses in the sequential loop. To counteract this, FLASH provides a mixed-chunk attention formulation, where the input sequence is chunked into $\lfloor L/C\rfloor$ chunks with fixed size $C$. From each chunk $c$, representations $R_{[c]},V_{[c]}\in\mathbb{R}^{C\times d_{e}}$ and $Z_{[c]}\in\mathbb{R}^{C\times d_{z}}$ are obtained following Eq. (26). Then, an attention mechanism composed by a local (quadratic) and a global (linear) component is formulated. Local attention follows the same procedure of Eq. (26) to compute a local chunk-based version of $\hat{V}$, denoted with $\hat{V}_{[c]}^{\text{loc}}$, given by $\hat{V}^{\text{loc}}_{[c]}=\text{ReLU}^{2}(\hat{Q}^{\text{loc}}_{[c]}\hat{K}^{{\text{loc}}^{\prime}}_{[c]})V_{[c]}$. Here $\hat{Q}^{\text{loc}}_{[c]}$ and $\hat{K}^{\text{loc}}_{[c]}$ are the outcome of two ad-hoc per-dim scaling/offsets transformations of $Z_{[c]}$. The complexity of these operations is linear in the sequence length, i.e., $\mathcal{O}(LC{d_{\text{in}}})$. Differently, the global linear attention captures long-interactions across chunks, exploiting other ad-hoc per-dim scaling and offsets transformations of $Z_{[c]}$, $\hat{Q}^{\text{glob}}_{[c]}$, and $\hat{K}^{\text{glob}}_{[c]}$. The causal formulation is defined as follows,
@@
[…]
@@
$\hskip-9.95863pt\text{Retention}(U)=(QK^{\prime}\odot D)V,$
@@
The intra-chunk component is a linear attention performed on the chunk, and thus takes $\mathcal{O}(C^{2}{d_{\text{k}}}+C{d_{\text{k}}}^{2})$, while the cross-chunk integrates the previous chunk component for the contribution from the hidden state from the previous chunk, and takes $\mathcal{O}(C{d_{\text{k}}}^{2})$. Overall, training complexity is $\mathcal{O}\left(\frac{L}{C}(C^{2}{d_{\text{k}}}+C{d_{\text{k}}}^{2})\right)=\mathcal{O}(LC{d_{\text{k}}}+L{d_{\text{k}}}^{2})$. The chunk size $C$ can be controlled for a trade-off between FLOPs and wall-clock speed. Overall, the decay factor introduced by RetNet puts more weight on recently processed inputs and is independent on the processed data. Finally, RetNet exploits multiple heads equipped with retention and a different $\gamma$ for each head, resulting in different variance statistics. GroupNorm [306] normalizes the output of each head, and a swish gate [103] increases the non-linearity of retention layers. RetNet is not the only approach belonging to this category. TransNormerLLM [231] is claimed to be the first linear attention-based Large Language Model (LLM) (up to 175 billion parameters) that outperforms conventional softmax attention-based models in terms of both accuracy and efficiency. It builds upon TransNormer[230] (see Section 3.1), replacing diagonal attention with linear attention. The model address the issue of attention dilution by adding linearized relative positional encodings with exponential decay [234], linear attention acceleration (by leveraging the recomputation technique from FlashAttention [54] to avoid the materialization of the 2D hidden state $S_{t}$ – see Section 8.2), tensor normalization from [230], and a gating mechanism with a decay factor applied to the additive recurrent update. GateLoop [135] incorporates a data-controlled gating mechanism which is applied on inputs, hidden states and outputs, replacing the fixed decay rate exploited in RetNet with a time-varying data-dependant diagonal state transition $A_{t}\in\mathbb{C}^{{d_{\text{s}}}\times{d_{\text{s}}}}$, defined in polar form,
$A_{t}$ $=\text{diag}(\gamma_{t}e^{i\theta_{t}})\vcentcolon=\text{diag}(\sigma(\alpha_{t})e^{i\beta_{t}}),$ (31)

### v2

@@ 3.3 Modeling Recurrence
where $\text{sg}(\cdot)$ denotes an operation (i.e., stop-gradient) which avoids the gradient flow during Backpropagation, $[\cdot|\cdot]$ denotes concatenation and $\text{tr}^{l}(\cdot)$ the $l$-th Transformer layer. Notice that the recurrent dependency shifts one-layer downward per segment, as depicted in Figure 8, which differs from the same-layer recurrence in conventional RNNs. Thus, the largest possible dependency length grows linearly w.r.t. the number of layers as well as the segment length. Additionally, it is possible to not limit the state-cache solely to the previous segment, by storing the last $m$ states and using them as the extra context when processing the current segment. Thus, a Transformer-XL with $N$ layers and with a memory of size $m$, has a maximum temporal range of $N\times m$ with a memory complexity in the attention operation of $\mathcal{O}(T^{2}+Tm)$ when processing a segment with length $T$.
@@
In RMT the Effective context length for recurrence with memory is not limited by the depth of the network, which is the case for the cache of Transformer-XL. Moreover, while Transformer-XL stores $m\times T$ vectors per segment, RMT stores only $m$ memory vectors per segment. RMT is trained with BPTT, and the number of segments to backpropagate is a hyperparameter of the training procedure (the authors tested from 0 to 4 segments). Differently from Transformer-XL, during the backward pass memory-related gradients are not stopped between segments. Figure 8 depicts the architectural differences between Transformer-XL and RMT. In a recent technical report [124], the authors leverage the RMT architecture to extend the context length of a BERT model [125] up to two million tokens, while maintaining high memory retrieval accuracy. Another approach that still falls in the category of models described so far is Infini-Transformer [126], which incorporates a compressive memory into a vanilla attention mechanism. It also builds in both causal local attention and long-term linear attention mechanisms into a single Transformer block. Inspired by Dai et al. [117], the authors remark that performing attention solely on the current segment (that can be seen as a form of local attention) discards the attention states of the previous one. To counteract this issue, they propose to maintain the entire context history in the form of a compressive memory. An Infini-Transformer layer contains both global compressive and local fine-grained states. In fact, it maintains as many parallel compressive memories as the number of attention heads, in addition to the dot-product attention. Each segment is processed via a classic softmax self-attention, producing an output $Y_{\text{loc}}$. Then, a linear attention (see Eq. (9)) is exploited to retrieve content from the previous segment memory/state, $M_{s-1}$, as follows,
@@
where $\phi(x)=\text{elu}(x)+1$, $s$ is an index over the segments and $z_{.}$ is the normalization factor of linear attention (see Eq. (9), here we used the parallel form, in matrix notation). State/memory update is implemented by a delta-rule-like mechanism [85] (see Section 3.1), first retrieving existing entries (values) and subtracting them from the new values, before applying the associative bindings, as follows,
@@
Chunk-level Recurrences.   The recently introduced segment-based approach allows networks to process very long texts sequentially, keeping a recurrent memory/context. Another category of emerging approaches involving recurrent models that handle portions of text consists of processing sub-portions of the input sequence, referred to as chunks, by dividing the sequence into non-overlapping chunks and performing (serial) inter-chunk recurrent computations followed by (parallel) intra-chunk computations. Such a chunk-wise parallel form yields sub-quadratic complexity. Temporal Latent Bottleneck (TLB) [127] divides the input sequence $U=(u_{1},\dots,u_{L})$ into chunks of fixed size $C$, resulting in $\lfloor L/C\rfloor$ chunks that are sequentially processed one after the other. We denote the $i$-th chunk of the input sequence as $U_{[i]}\vcentcolon=U_{iC:(i+1)C}\in\mathbb{R}^{C\times{d_{\text{in}}}}$. Each chunk is processed by a fast-stream (also referred to as perceptual module), implemented by a Transformer $\text{tr}(\cdot)$. The fast-stream computation is conditioned via cross-attention on information coming from a slow-stream, referred to as Temporal Latent Bottleneck $\mathcal{G}$, that aggregates cues across chunks and is updated once per chunk. Such a slow-stream is computed recurrently, and it manages a state X{X} composed of a set of NN $d_{\text{s}}$}-dimensional vectors. The state update in $\mathcal{G}(\cdot)$ is performed via a cross-attention operation in which the queries are obtained by projecting $X$, while the keys and values come from the output of the perceptual module. Overall, the model performs the following operation,
@@
The recurrent update of the TLB state $X$ is performed at lower rates with respect to the computations happening in the perceptual module, fostering the distillation of more stable, condensed and slowly changing features, while the perceptual module captures faster changing local cues. From the computational point of view, leveraging a chunk-based approach allows to achieve a complexity of $\mathcal{O}\left(\frac{L}{C}(C^{2}d+CN)\right)$, where $N$ is the number of temporal latent bottleneck state vectors. Hence, it has a much lower computational complexity compared to a vanila Transformer applied on the overall sequence. A concurrent work, Block-Recurrent Transformer (Brect) [128], handles the input sequence in a similar manner to Transformer-XL: a document is split into multiple segments of size $T$, processed one after the other. Each segment is processed in chunks (or blocks) using a sliding window attention with size $C$, with a causal mask that forces each token to attend solely to the previous ones. Brect is composed of a recurrent cell that receives as inputs $C$ token embeddings, where $C$ is the block/window size, and a set of $N$ state vectors. Similarly to TLB, a two-way processing is performed in the proposed recurrent cell, consisting of the so-called vertical and horizontal “directions”: the vertical direction (i.e., corresponding to the fast-stream in TLB) is implemented by a Transformer layer that performs self-attention over the input tokens of the current block and cross-attends to the recurrent states, producing output token embeddings for the current block. The horizontal direction (i.e., the slow-stream in TLB) performs self-attention over the current state vectors, and cross-attends to the block input tokens, producing the next state vectors. Unlike TLB, cross-attention and self-attention are computed in parallel. Recurrence is integrated with the sliding window attention mechanism, since keys and values from the previous window are stored in a differentiable cache and concatenated to the ones of the current block. Residual connections are replaced by gating mechanisms (implemented either as fixed convex combinations or trainable LSTM-like gates) that help the model forget redundant information. For every layer of the architecture, the last recurrent states of a segment are stored in a non-differentiable cache and fed to the following segment in the document as a warm-start. This mechanism extends the sliding window to cover the entire sequence. The cache implements a form of truncated BPTT over long documents. Block-State Transformer [129] replaces the recurrent cell in Brect with a linear state-space model (see Section 4), which processes the sequence and preserves long-range dependencies, while allowing also for parallel computations.
Recurrent Attention Networks (RAN)[130] iteratively process the input sequence by means of non-overlapping windows, each of them processed via multi-head self-attention. Intra-window information is propagated by a global perception cell (GPC) vector, extracted from the self-attention representation of the current window. The GPC vector is concatenated with the input tokens of the next window. A memory review mechanism cross-attends the concatenation of the GPC vectors of all the windows to build the final output, encoding both contextual and global information. We report in Figure 9 the main differences in the processing schemes of segment level and chunk-level approaches, reporting examples of some of the described models.
@@ 3.4 Decaying Linear Attention by Gating Functions
Gating mechanisms were introduced since the dawn of RNNs, being a key feature of the popular LSTM model [28, 131]. In previous subsections and in the context of linear approximations to attention mechanisms, we described several works connecting Linear Transformers to Fast weight Programmers (FWP), which introduced decay/gating mechanisms inherited from FWP [85, 94, 96, 98]. The original gating functions in LSTMs consisted of units with sigmoidal activations, responsible for gating specific signals (input, output, state) via multiplicative operations. Lately, the concept of gating has been relaxed to consider any multiplicative interaction, possibly followed by an activation function (e.g., element-wise multiplicative components that do not interact along the sequence length are referred to as gating mechanisms [132, 133]). Nevertheless, the role of gating has become increasingly popular and now pivotal in many works.
Multiplicative Interactions.   Hua et al. [132] remarked that despite the linear theoretical complexity claimed by linear attention methods [72], they have not been able to supersede vanilla Transformers as the dominant choice in state-of-the-art systems. They attributed this to ($i$) approximations needed to achieve efficiency; ($ii$) non-trivial gap between theoretical complexity and empirical speed on accelerators, due to memory re-formatting and operations (see also Section 8.2); (iiiiii) slow training on causal autoregressive tasks, due to sequential processing of the recurrent forms. To counteract these issues, the authors proposed FLASH, which is based on a novel layer, dubbed Gated Attention Unit (GAU), to substitute softmax multi-head self-attention. The key idea is to combine a Gated Linear Unit (GLU) [^10] A GLU [134] is an MLP which output is modulated via a gating (i.e., a learned multiplicative interaction). In fact, the layer input $U$ is projected by two learnable matrices $W_{p}\in\mathbb{R}^{{d_{\text{in}}}\times d_{e}}$ and $W_{t}\in\mathbb{R}^{{d_{\text{in}}}\times d_{e}}$ into two representations $P,T\in\mathbb{R}^{L\times d_{e}}$, that interact in an element-wise multiplicative manner to produce the layer output, as follows: $R=\sigma(UW_{p})\quad V=\sigma(UW_{t})\quad Y=(R\odot V)W_{y}.$ [134] and attention in a unified layer, as depicted in Figure 10.
@@
where $W_{r},W_{v}\in\mathbb{R}^{{d_{\text{in}}}\times d_{e}}$, $W_{z}\in\mathbb{R}^{{d_{\text{in}}}\times d_{z}}$ with $d_{z}\ll{d_{\text{k}}}$, $\mathcal{Q},\mathcal{K}$ are learned transformations that apply per-dim scalars and offsets to $Z$, $W_{y}\in\mathbb{R}^{d_{e}\times{d_{y}}}$ denotes an output learnable matrix, and $b_{v}\in\mathbb{R}^{d_{z}}$ are relative position bias. Notice that this formulation substitutes the softmax with the squared ReLU[107] (denoted with $\text{ReLU}^{2}$). This single-headed softmax-free formulation is capable to achieve multi-head Transformers performances without quality loss, and is cheaper in terms of computational requirements, by adding solely the learnable matrix $W_{z}$ with ${d_{\text{in}}}\times z$ parameters on top of the GLU. Additionally, the authors analyze the causes underlying the aforementioned memory and speed issues (issues $ii$ and iiiiii). In fact, despite the huge advantages brought by the constant-inference computation and memory cost in autoregressive decoding (i.e., thanks to the stateful representation $S_{t}$) the rearrangement of computations in linear attention lead to an inefficiency in the case of autoregressive training. [^11] Due to the causal constraint for auto-regressive training, the query vector corresponds to a different cache value at each time step. This requires the model to compute and cache $L$ different values of the incremental state and requires $L$ memory accesses in the sequential loop. To counteract this, FLASH provides a mixed-chunk attention formulation, where the input sequence is chunked into $\lfloor L/C\rfloor$ chunks with fixed size $C$. From each chunk $c$, representations $R_{[c]},V_{[c]}\in\mathbb{R}^{C\times d_{e}}$ and $Z_{[c]}\in\mathbb{R}^{C\times d_{z}}$ are obtained following Eq. (26). Then, an attention mechanism composed by a local (quadratic) and a global (linear) component is formulated. Local attention follows the same procedure of Eq. (26) to compute a local chunk-based version of $\hat{V}$, denoted with $\hat{V}_{[c]}^{\text{loc}}$, given by $\hat{V}^{\text{loc}}_{[c]}=\text{ReLU}^{2}(\hat{Q}^{\text{loc}}_{[c]}\hat{K}^{{\text{loc}}^{\prime}}_{[c]})V_{[c]}$. Here $\hat{Q}^{\text{loc}}_{[c]}$ and $\hat{K}^{\text{loc}}_{[c]}$ are the outcome of two ad-hoc per-dim scaling/offsets transformations of $Z_{[c]}$. The complexity of these operations is linear in the sequence length, i.e., $\mathcal{O}(LC{d_{\text{in}}})$. Differently, the global linear attention captures long-interactions across chunks, exploiting other ad-hoc per-dim scaling and offsets transformations of $Z_{[c]}$, $\hat{Q}^{\text{glob}}_{[c]}$, and $\hat{K}^{\text{glob}}_{[c]}$. The causal formulation is defined as follows,
@@
where $\dagger$} denotes the conjugate transpose. Notice the resemblance between the multiplying factors and the xPOS [138] positional encodings. Starting from this formulation, it is easy to obtain the RetNet parallel form (Figure 11, left), which is defined as follows when considering a vector mapping instead of the scalar one of Eq. (29),
@@
$\thinspace\text{Retention}(U)=(QK^{\prime}\odot D)V,$
@@
The intra-chunk component is a linear attention performed on the chunk, and thus takes $\mathcal{O}(C^{2}{d_{\text{k}}}+C{d_{\text{k}}}^{2})$, while the cross-chunk integrates the previous chunk component for the contribution from the hidden state from the previous chunk, and takes $\mathcal{O}(C{d_{\text{k}}}^{2})$. Overall, training complexity is $\mathcal{O}\left(\frac{L}{C}(C^{2}{d_{\text{k}}}+C{d_{\text{k}}}^{2})\right)=\mathcal{O}(LC{d_{\text{k}}}+L{d_{\text{k}}}^{2})$. The chunk size $C$ can be controlled for a trade-off between FLOPs and wall-clock speed. Overall, the decay factor introduced by RetNet puts more weight on recently processed inputs and is independent on the processed data. Finally, RetNet exploits multiple heads equipped with retention and a different $\gamma$ for each head, resulting in different variance statistics. GroupNorm [139] normalizes the output of each head, and a swish gate [140] increases the non-linearity of retention layers. RetNet is not the only approach belonging to this category. TransNormerLLM [141] is claimed to be the first linear attention-based Large Language Model (LLM) (up to 175 billion parameters) that outperforms conventional softmax attention-based models in terms of both accuracy and efficiency. It builds upon TransNormer[103] (see Section 3.1), replacing diagonal attention with linear attention. addresses the issue of attention dilution by adding linearized relative positional encodings with exponential decay [142], linear attention acceleration (by leveraging the recomputation technique from FlashAttention [37] to avoid the materialization of the 2D hidden state $S_{t}$ – see Section 8.2), tensor normalization from [103], and a gating mechanism with a decay factor applied to the additive recurrent update. GateLoop [143] incorporates a data-controlled gating mechanism which is applied on inputs, hidden states and outputs, replacing the fixed decay rate exploited in RetNet with a time-varying data-dependant diagonal state transition $A_{t}\in\mathbb{C}^{{d_{\text{s}}}\times{d_{\text{s}}}}$, defined in polar form,
$A_{t}$ $=\text{diag}(\gamma_{t}e^{i\theta_{t}})=\text{diag}(\sigma(\alpha_{t})e^{i\beta_{t}}),$ (31)
@@
This section shows that the once-sharp boundary between self-attention and recurrence is rapidly eroding. Modern Transformer research first adopts kernel or linear attention schemes (e.g., Linear Transformer, Performer) that rewrite the softmax dot product as a feature-kernel product. This reframes attention as an additive update of a constant-size matrix state and cuts the quadratic $\mathcal{O}(L^{2})$ cost of vanilla attention to linear $\mathcal{O}(L)$ while preserving full parallel training. Extending that idea, fast-weight and delta-rule methods such as DeltaNet and SRWM treat the key–value cache as a programmable associative memory, using outer-product updates and data-controlled forgetting gates to balance capacity and recency. A parallel line of work eliminates dot-products entirely: low-rank or channel-wise interactions in Attention-Free Transformers and the RWKV family compute gated exponential averages that naturally down-weight distant tokens and enable linear-time decoding. Finally, segment and chunk recurrent architectures such as Transformer-XL, RetNet and their variants propagate compressed hidden states across windows, extending effective context length without incurring quadratic growth. Together these threads reveal a unifying trend: contemporary Transformers increasingly resemble specialised recurrent networks, compressing history into compact learned states, applying explicit decay or memory tokens to regulate long-range influence, and thereby reclaiming much of the efficiency once unique to classic RNNs.
While this section demonstrated that many Transformer variants now reuse recurrent ideas, an orthogonal line of work pursues the same goal from first principles: State-Space Models (SSMs). By discretising continuous-time dynamics, SSMs encode sequence history in compact linear states and thus inherit the favourable $\mathcal{O}(L)$ scaling of classical RNNs without relying on attention. Section 4 reviews this rapidly growing family and its role in long-sequence processing.